OpenStudy (anonymous):

How do you find the cube root of -1+isqrt3?!?!?!

OpenStudy (anonymous):

Is this the equation?\[\sqrt[3]{-1+i}\] or is it \[-1+i \sqrt{3}\]?

OpenStudy (anonymous):

it is the bottom one

OpenStudy (anonymous):

@Johnbc

geerky42 (geerky42):

So you need to find cube root of bottom one Johnbc gave you?

OpenStudy (anonymous):

yes

geerky42 (geerky42):

You would need to convert −1+i√3 to cosθ + i sin θ form

geerky42 (geerky42):

rcosθ + i r sinθ* You then would need to apply DeMoivre's Theorem \(\Large [r(\cos\theta + i \sin\theta)]^n = r^n(\cos(n\theta) + i\sin(n\theta))\) Do you know how to convert it to r (cosθ + i sinθ) form?

OpenStudy (anonymous):

no i dont.... i tried but i think i messed it up a lot. @geerky42

geerky42 (geerky42):

hmm, we would need r and θ first. Do you know how to find r?

OpenStudy (anonymous):

i think its \[r|z|=\sqrt(a^2+b^2)\]

OpenStudy (anonymous):

@geerky42

OpenStudy (anonymous):

i tried that but oh well ill show you: \[-1+isqrt3\] \[z=r(\cos \Theta+isin \Theta)\] \[r|z|=\sqrt(a^2+b^2)\] \[\sqrt(-1^2+(\sqrt3^2))\] \[=\sqrt2\]

OpenStudy (anonymous):

@geerky42